# Learning Satellite Attitude Dynamics with Physics-Informed Normalising Flow

### v1

@@ 3.1 Satellite Attitude Dynamics
$\Omega[\omega]=\begin{bmatrix}q_{0}&-q_{3}&q_{2}\\ q_{3}&q_{0}&-q_{1}\\ -q_{2}&q_{1}&q_{0}\end{bmatrix}$ (6)
@@ 3.4 Physics-Informed Training
Here $\alpha$ and $\beta$ are used to change the relative weight of the two loss functions. We experimented with fixed hand-selected values and with the Lagrangian dual approach (Fioretto et al. (2021)), which automatically learns the best multipliers with a sub-gradient method during training. In the latter case $\alpha=(1-\beta)$, and $\beta$ was constrained in the interval [0, 1].
@@ 4.1 Dataset
We ran 300 runs randomly varying the initial attitude to generate an initial dataset, and other 50 runs with a different satellite inertia matrix and mass, to test the robustness of the trained models to errors of approximately 10% on all axes, as these parameters are typically not known with absolute precision and may change during the operational life. In each simulation, the satellite started with a random angular velocity of the reaction wheels, sampled uniformly from [-300, 300] rotations per minute. Each simulation lasted 3 minutes with a sampling and control time of 0.1 seconds, and a simulation time step of 0.001 seconds with Runge-Kutta 4 as integrator.
We construct the dataset by creating input-output pairs as follows: the input $x_{i}$ has shape (1, 12) and is defined as $x_{i}=\{\omega,\omega_{rw},u_{rw},\dot{\omega}\}$, while the target $y_{i}$ is equal to the change in satellite angular velocity, i.e. $y_{i}=\Delta\omega_{i}$.

### v2

@@ 3.1 Satellite Attitude Dynamics
$\Omega[\omega]=\begin{bmatrix}0&-\omega_{0}&-\omega_{1}&-\omega_{2}\\ \omega_{0}&0&\omega_{2}&-\omega_{1}\\ \omega_{1}&-\omega_{2}&0&\omega_{0}\\ \omega_{2}&\omega_{1}&-\omega_{0}&0\end{bmatrix}$ (6)
@@ 3.4 Physics-Informed Training
Here $\alpha$ and $\beta$ are used to change the relative weight of the two loss functions. We experimented with fixed hand-selected values and with the Lagrangian dual approach (Fioretto et al. (2021)), which automatically learns the best multipliers with a sub-gradient method during training, dynamically changing the relative importance of the two losses based on the performance of the model. In the latter case $\alpha=(1-\beta)$, and $\beta$ was constrained in the interval [0, 1].
@@ 4.1 Dataset
The rationale for this dataset design is to reproduce a representative operational environment for small satellites performing attitude control solely with RWs. By including both nominal and perturbed configurations, the dataset allows assessing not only nominal performance but also the robustness of the learned models to model uncertainties and off-nominal conditions. We ran 300 runs randomly varying the initial attitude, and orbital position, to generate an initial dataset, and other 50 runs with a different satellite inertia matrix and mass, to test the robustness of the trained models to errors of approximately 10% on all axes, as these parameters are typically not known with absolute precision and may change during the operational life. These trajectories cover a broad portion of the attitude space and reaction wheel state space, providing a diverse training and testing distribution of control inputs and system responses. In each simulation, the satellite started with a random angular velocity of the reaction wheels, sampled uniformly from [-300, 300] rotations per minute. Each simulation lasted 3 minutes with a sampling and control time of 0.1 seconds, and a simulation time step of 0.001 seconds with Runge-Kutta 4 as integrator.
We construct the dataset by creating input-output pairs as follows: the input $x_{i}$ has shape (1, 12) and is defined as $x_{i}=\{\omega,\omega_{rw},u_{rw},\dot{\omega}\}$, while the target $y_{i}$ is equal to the change in satellite angular velocity, i.e. $y_{i}=\Delta\omega_{i}$. The training dataset was split 67-33% respectively for training and validation.